OpenStudy (anonymous):

the sum of two numbers is 13/6.An even number of arithematic means are being inserted between them and their sum exceeds their number by 1.Find the number of means inserted.

OpenStudy (anonymous):

I'm not sure I understand the question...This is confusing and ambiguous.

OpenStudy (anonymous):

say the numbers are a and b. The arithmetic mean of a and b is (a+b)/2. so the arithmetic mean of these two numbers is 13/12, because they already gave us a+b=13/6. now we're inserting a number (n) of these, and their sum exceeds their number by 1: n(13/12) = n + 1 13n/12 = n + 1 13n = 12n + 12 n = 12